OpenStudy (anonymous):

How many moles of N are in 0.203 g of N2O?

OpenStudy (anonymous):

0.203 g N2O x \[\frac{ 1 mol }{ 44.014 g N2O }\] x \[\frac{ 1 mol N2O }{ 2 mol N }\] = .0023 moles of N

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

oops wait I think the final fraction might need to be flipped

OpenStudy (anonymous):

the final amount cant be less..

OpenStudy (anonymous):

Okay so multiply \[0.203 g N _{2}O * \frac{ 1 mol N _{2}O }{ 44.014 g N _{2}O } * \frac{ 2 mol N }{ 1 mol N _{2}O } = .0092 mol\]

OpenStudy (anonymous):

You just have to do this like the last one i showed you!